Assume that over a given period the price index (P) is 5, the volume of the transactions (T) is 2,000 and money supply (M) is 1,000. What is the velocity of money in this economy?

a.

10

b.

6

c.

4

d.

8.

2.Which macroeconomic model is defined below?

The productivity of labor depends on the real wage workers are paid. In such models, the real wage is set to maximize the productivity units of labor per dollar of expenditure, not to clear the labor market. For this reason, we observe on the market higher wages than its equilibrium level. This causes the involuntary unemployment problem.

a.

Simple Real Business Cycle (RBC) model

b.

Insider-outsider model

c.

Menu cost model

d.

Efficiency wage model


1
Expert's answer
2021-07-12T07:30:02-0400

1. Velocity of money is given by the product of price index (P) and volume of transactions (T) divede by money supply.

"Velocity of Money =\frac {P\times T}{M}"

"Velocity of money = \frac {5 \times 200}{1000}"

"=10"

The answer is a. 10


2. Answer is d. Efficiency wage model.